1.   The sequence {Un} is defined by the first term u1=-2 and the recurrent relation : Un+1=Un/4 +3

a.   Evaluate U2,U3,U4

b.    Consider  the sequence Vn such that Vn=Un+1-Un for n>=2

c.    Prove that {Vn} is a G.P 

d.   Express Vn  and Un  in terms of n.

e.     Evaluate the limits of  Un and Vn  

f.    Is Un an increasing  or decreasing sequence ? 


1
Expert's answer
2022-05-02T15:17:35-0400

a.

"u_1=-2,"

"u_2=-2/4+3=\dfrac{5}{2},"

"u_3=(\dfrac{5}{2})/4+3=\dfrac{29}{8},"

"u_4=(\dfrac{29}{8})/4+3=\dfrac{125}{32}"

b.


"v_n=u_{n+1}-u_n, n\ge2"

"v_n=u_n/4+3-u_n, n\ge2"

"v_n=3-3u_n/4, n\ge2"

c.

"v_{n+1}=3-3u_{n+1}\/4=3-3(u_n\/4 +3)\/4="

"=(3-3u_n/4)/4=v_n/4, n\ge2"

"\{v_n\}" is a G.P 

"v_2=\dfrac{9}{8}, q=\dfrac{1}{4}"


d.


"v_n=\dfrac{9}{8}(\dfrac{1}{4})^{n-2}, n\ge2"

"u_n=4-\dfrac{4v_n}{3}, n\ge2"

"u_n=4-\dfrac{3}{2}(\dfrac{1}{4})^{n-2}, n\ge2"

"u_1=-2"

e.


"\lim\limits_{n\to\infin}u_n=\lim\limits_{n\to\infin}(4-\dfrac{3}{2}(\dfrac{1}{4})^{n-2})=4"

"\lim\limits_{n\to\infin}v_n=\lim\limits_{n\to\infin}(\dfrac{9}{8}(\dfrac{1}{4})^{n-2})=0"

f.

"\{u_n\}" is increasing sequence.